# A priori error analysis of the proximal Galerkin method

### v1

@@ 5. Application II: The Signorini problem
Here, the boundary of $\Omega$ consists of two (relatively) open, disjoint subsets $\partial\Omega=\overline{\Gamma_{\rm T}\cup\Gamma_{\rm D}}$ with $|\Gamma_{\rm T}|>0$ and $|\Gamma_{\rm D}|>0$. For further simplicity, we assume that the contact boundary $\Gamma_{\rm T}$ is an open straight line segment and fix a gap function $g\in H^{1}(\Omega)$ with $g|_{\Gamma_{\rm D}}=\delta>0$. Recall that in this case, $\lambda^{*}=(\mathsf{C}\epsilon(u^{*})n)\cdot n$.
@@
(136b) $W_{h}$ =$=\{w_{h}\in C(\overline{\Gamma_{\mathrm{T}}})\mid w_{h}|_{E}\in\mathbb{P}_{1}(E)~{}\text{ for all }\text{edges }E\subset\Gamma_{\mathrm{T}},\;\;w_{h}=0\text{ on }\partial\Gamma_{\mathrm{T}}\}.$
@@
We provide Algorithm 3, the application of the PG method (Algorithm 1) to this problem for completeness; see also [21, Example 2].
@@
[…]
@@
In addition, the Fortin operator $\Pi_{h}:H^{1}_{\mathrm{D}}(\Omega)^{2}\rightarrow V_{h}$ given by $\Pi_{h}w=(\tilde{\Pi}_{h}w_{n})n+(\mathcal{SZ}_{h}w_{\tau})\tau$ satisfies 42 and the stability and approximation bounds 83a, 83b and 85a stated in Assumption 3.
@@
Combining 143 and 142 shows that 83a holds.
We proceed to verify the error estimate 85a. Note that $\Pi_{h}v_{h}=v_{h}$ for any $v_{h}\in\mathbb{P}_{1}(\mathcal{T}_{h})\cap H^{1}_{\rm D}(\Omega)$. To see this, note that $\pi_{h}v_{h}=v_{h}|_{\Gamma_{\rm{T}}}$ since $v_{h}|_{\Gamma_{\mathrm{T}}}\in W_{h}$ and $\mathcal{SZ}_{h}v_{h}(z)=v_{h}(z)$ for all $z\in\mathcal{N}_{h}$ [46]. Hence,
@@
where we used the just proven stability property 83a of $\Pi_{h}$. The estimate on the $L^{2}$ error stated in 85a can now be concluded from the approximation properties of $\mathcal{SZ}_{h}$. To show 83b, one uses an inverse estimate followed by approximation properties. Namely, we have
@@
where we used the stability of $\mathcal{SZ}_{h}$ in the $H^{1}(\Omega)$-seminorm, see [9, Corollary 4.8.15]. The final error estimate for $\tilde{\Pi}_{h}$ in the $H^{1}(\Omega)$-seminorm is proven similarly. ∎
In order to check Assumption 2, we observe that the set $K_{h}$ 62 satisfies
@@
We now define ${\alpha}_{z,i}$. Let $(\hat{w}_{j},\hat{x}_{j})$, $j=0,1$, be the non-negative weights and points given by the Gauss–Radau quadrature rule [25, Table 6.1], which is an exact quadrature rule for polynomials of degree $2$ on the reference facet $\hat{E}$. Let $F_{E}$ denote the affine linear transformation from the reference facet $\hat{E}$ to a facet $E\subset\Gamma_{\mathrm{T}}$. Observe that these quadrature points define weights and points $(w_{j},x_{j})$ with $x_{j}\in\tilde{\omega}_{z}$ for $j=0,1,2,3$, such that for any continuous piecewise second-order polynomial $v$ on $\tilde{\omega}_{z}=E_{1}\cup E_{2}$, we can write
@@
Further, Assumption 5 guarantees that the weights corresponding to the boundary nodes in $\partial\Gamma_{\mathrm{T}}$ are zero in 150; i.e., $\alpha_{z}=0$ if $z\in\partial\Gamma_{\mathrm{T}}$. In turn, for each $z\in\mathcal{N}_{h}\cap\Gamma_{\mathrm{T}}$, there exists an element $T_{z}\in\mathcal{T}_{h}$, $T_{z}\subset\cup_{z_{i}}\omega_{z_{i}}$, such that if $p\in\mathbb{P}_{1}(\Delta_{T_{z}})$, it holds that
@@
[…]
@@
We first consider the error $(\mathcal{C}_{h}v_{h}-v_{h})$ for any $v_{h}\in V_{h}$. For a node $z\in\Gamma_{\mathrm{T}}$, we consider the element $T_{z}$ such $\mathcal{C}_{h}p(z)=p(z)$ for $p\in\mathbb{P}_{1}(\Delta_{T_{z}})$; cf. 151. Then,
@@
Using that $\|\varphi_{z}\|_{L^{2}(T)}\lesssim h_{T}$, 155, and applying Cauchy–Schwarz inequality, we obtain that
@@
Let $u^{*}$ solve 134 and let $(u_{h}^{\ell},\psi_{h}^{\ell})$ come from Algorithm 3. Assume that $u^{*}\in H^{1+s}(\Omega)^{2}$, $g\in H^{1+s}(\Omega)$ and $-\mathrm{div}(\mathsf{C}\epsilon(u))-f\in H^{-1+r}(\Omega)^{2}$ for some $r,s\in(0,1]$, $\|\psi_{h}^{0}\|_{L^{\infty}(\Gamma_{\rm T})}\lesssim 1$, and Assumption 5 holds. The following error estimate holds:
@@
(160) $(\mathcal{E}_{h}w)_{n}=\mathcal{C}_{h}(w_{n}-g)+g-\tilde{\epsilon}_{h},\quad(\mathcal{E}_{h}w)_{\tau}=\mathcal{SZ}_{h}w_{\tau},$
where
$\tilde{\epsilon}_{h}=\sum_{z\in\mathcal{N}_{h}\cap\Gamma_{\mathrm{T}}}\epsilon\varphi_{z},\quad\epsilon=\mathrm{card}(\mathcal{T}_{h})^{-1/2}\min_{T\in\mathcal{T}_{h}}h_{T}.$
To check Assumption 2, we first note that $(\mathcal{E}_{h}w)_{n}=0$ on $\Gamma_{\rm D}$, $(\mathcal{E}_{h}w)_{n}\in H_{\mathrm{D}}^{1}(\Omega)$ and $(\mathcal{E}_{h}w)_{\tau}\in H_{\mathrm{D}}^{1}(\Omega)$ from the properties of $\mathcal{SZ}_{h}$. This implies that $\mathcal{E}_{h}w\in H^{1}_{\mathrm{D}}(\Omega)^{2}$. We now check that $\operatorname{ess\,inf}(g-\mathcal{E}_{h}w\cdot n)>0$ in $\Gamma_{\mathrm{T}}$ for any $w\in K_{h}\cup K$ which implies that $\mathcal{E}_{h}w\in\mathrm{dom}(\nabla\mathcal{R})$, where $\nabla\mathcal{R}(u)=\ln(g-u)$, see 137.
For $w\in K\cup K_{h}$, we have that $(u_{h}\cdot n-g,\varphi_{z})_{\tilde{\omega}_{z}}\leq 0$ for any $z\in\Gamma_{\mathrm{T}}$ since the support of $\varphi_{z}$ on $\Gamma_{\mathrm{D}}$ is $\tilde{\omega}_{z}$. The additional Assumption 5 on the mesh guarantees that $\tilde{\omega}_{z}$ for $z\in\partial\Gamma_{\rm T}$ is not included in the definition of $\mathcal{C}_{h}$ 147. From 147, we now obtain that $\mathcal{C}_{h}(w_{n}-g)(z)\leq 0$ for all $z\in\Gamma_{\mathrm{T}}$. In addition, for the nodes on $\partial\Gamma_{\rm T}$, we find $\mathcal{C}_{h}(w_{n}-g)=-g|_{\Gamma_{\rm D}}=-\delta<0$. Hence, on $\Gamma_{\rm T}$, we conclude that
$\mathcal{C}_{h}(w_{n}-g)-\tilde{\epsilon}_{h}\leq-\sum_{z\in\partial\Gamma_{\rm T}}\delta\varphi_{z}-\sum_{z\in\mathcal{N}_{h}\cap\Gamma_{\mathrm{T}}}\epsilon\varphi_{z}\leq-\min(\delta,\epsilon)\sum_{z\in\mathcal{N}_{h}\cap\overline{\mathrm{\Gamma}_{T}}}\varphi_{z}=-\min(\delta,\epsilon).$
Since $g-\mathcal{E}_{h}w\cdot n=\tilde{\epsilon}_{h}-\mathcal{C}_{h}(w_{n}-g)$, we conclude that $\operatorname{ess\,inf}(g-\mathcal{E}_{h}w\cdot n)\geq\min(\delta,\epsilon)>0$. This verifies Assumption 2.
Noting that $\mathcal{E}_{h}w=(\mathcal{C}_{h}w_{n},\mathcal{SZ}_{h}w_{\tau})+(g-\mathcal{C}_{h}g-\tilde{\epsilon}_{h},0)$, Assumption 3 is verified by applying Lemma 5.1, Lemma 5.2, standard estimates on $\varepsilon=(g-\mathcal{C}_{h}g-\tilde{\epsilon}_{h},0)$ similar to 132, and the approximation properties of the Scott–Zhang interpolant [46]. Details are skipped for brevity. ∎
@@ Acknowledgments
We thank Ioannis P. A. Papadopoulos for his comments on the manuscript and Thomas M. Surowiec and Dohyun Kim for numerous interesting discussions. BK and RM were supported in part by the U.S. Department of Energy Office of Science, Early Career Research Program under Award Number DE-SC0024335. Keith’s work was also supported in part by the Alfred P. Sloan Foundation via a Sloan Research Fellowship in Mathematics. MZ acknowledges support from an ETH Postdoctoral Fellowship for the project “Reliable, Efficient, and Scalable Methods for Scientific Machine Learning”.
@@ Appendix A A stability result
The generalized Shannon entropy 30 does not fit into the form 161 for all obstacles $\phi\in L^{\infty}(\Omega_{d};\mathbb{R}^{m})$. However, the decomposition 161 is written with sufficient generality to accommodate the settings analyzed in this paper. In particular, consider the obstacle problem from Section 4, where $\phi\in H^{1}(\Omega)$ and $\phi|_{\partial\Omega}=-\delta$ for a positive constant $\delta$. In this case, 161 holds with $\mathcal{R}_{0}^{*}(\psi)=\exp(\psi)-\delta\psi$, $\phi_{0}=1$, and $\phi_{1}=\phi+\delta$, implying $\nu_{1}=\delta$. For the Fermi–Dirac entropy given in Example 2.2, which is a suitable choice for bilateral obstacle problems ($B=\mathrm{id}$ and $Q=V=H^{1}_{0}(\Omega)$), we can write
@@
Thus, defining $\mathcal{R}_{0}^{*}(\psi)=2\ln(\exp(\psi)+1)-\psi$ and assuming that $\overline{u}-\underline{u}\geq\underline{\phi_{0}}$ and $\overline{u}+\underline{u}\in H^{1}_{0}(\Omega)$, we deduce that 161 holds with $\nu_{1}=1$. Notably, if $\underline{u},\overline{u}\in\mathbb{R}$ with $\underline{u}<0<\overline{u}$, the condition $\underline{u}+\overline{u}\in H^{1}_{0}(\Omega)$ is not generally satisfied. However, in this case, we can still verify 161 using an alternative decomposition given by $\mathcal{R}_{0}^{*}(\psi)=(\overline{u}-\underline{u})\ln(\exp(\psi)+1)+\underline{u}\psi$, $\phi_{0}=1$, and $\phi_{1}=0$. The Hellinger entropy, introduced in Example 2.3, readily satisfies 161 with $\mathcal{R}_{0}^{*}(\psi)=\gamma\sqrt{1+|\psi|^{2}}$, $\phi_{0}=1$, and $\phi_{1}=0$, implying $\nu_{1}=\gamma$ and $c_{1}=0$.
@@
Assume that 161 holds. Further, assume that $\psi_{h}^{0}$ is selected so that $\|B^{\prime}\psi_{h}^{0}\|_{V^{\prime}}\lesssim 1$. Then there exists a constant $C_{\mathrm{stab}}$, independent of $h$, $\ell$, and $\{\alpha_{k}\}_{k=2,\ldots}$, such that
@@
We first observe that testing 10a with $v_{h}=u_{h}^{k}$ and 10b with $w_{h}=\psi_{h}^{k}$, subtracting the resulting equations, and using the expression 161a yields
@@
Noting that $c_{1}\leq\mathcal{R}_{0}^{*}(0)$ by 161b, the above estimates yield
@@
Employing the coercivity of the bilinear form $a$ 7, we obtain that
@@
where $\beta>0$ is the same constant appearing in the LBB condition 41. Using the above bound in 167 and appropriate applications of Young’s inequality shows that there exists a constant $M_{k}$ (depending only on $\alpha_{k}$, $M$, $\|F\|_{V^{\prime}}$, $\|\Pi_{h}\|$, $\|\phi_{1}\|_{Q}$, $\beta$, and $c_{2}$) such that
@@
Using 7, we obtain that
@@
where we used (10a) and similar arguments to (169). The bounds 174 and 175 yield the first two terms in 163.
@@
where $C_{3}=(\|F\|_{V^{\prime}}+2\nu^{-1/2}MC_{2}+\|B^{\prime}\psi_{h}^{0}\|_{V^{\prime}})\|\Pi_{h}\|$. To derive the bound on $\|\psi_{h}^{k}\|_{L^{1}(\Omega_{d})}$, we substitute 178 into 167. ∎

### v2

@@ 5. Application II: The Signorini problem
Here, the boundary of $\Omega$ consists of two (relatively) open, disjoint subsets $\partial\Omega=\overline{\Gamma_{\rm T}\cup\Gamma_{\rm D}}$ with $|\Gamma_{\rm T}|>0$ and $|\Gamma_{\rm D}|>0$. For further simplicity, we assume that the contact boundary $\Gamma_{\rm T}$ is an open straight line segment and fix a gap function $g\in H^{1}(\Omega)$ with $g|_{\Gamma_{\rm D}}=\delta{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\geq}0$. Recall that in this case, $\lambda^{*}=(\mathsf{C}\epsilon(u^{*})n)\cdot n$.
@@
(126b) $W_{h}$ $=\{w_{h}\in C(\overline{\Gamma_{\mathrm{T}}})\mid w_{h}|_{E}\in\mathbb{P}_{1}(E)~\text{ for all }\text{edges }E\subset\Gamma_{\mathrm{T}},\;\;w_{h}=0\text{ on }\partial\Gamma_{\mathrm{T}}\}.$
@@
We provide Algorithm 3, the application of the PG method (Algorithm 1) to this problem for completeness; see also [19, Example 2].
@@
(129a) $\alpha_{k}\,(\mathsf{C}\,\epsilon(u^{k}_{h}),\epsilon(v_{h}))+(v_{h}\cdot n,\psi^{k}_{h}-\psi_{h}^{k-1})_{\Gamma_{\mathrm{T}}}$ $=\alpha_{k}\,(f,v_{h})$  $~\text{ for all }v_{h}\in V_{h},$
(129b) $(u^{k}_{h}\cdot n,w_{h})_{\Gamma_{\mathrm{T}}}+(\exp(-\psi^{k}_{h}),w_{h})_{\Gamma_{\mathrm{T}}}$ $=(g,w_{h})_{\Gamma_{\mathrm{T}}}$  $~\text{ for all }w_{h}\in W_{h}.$
@@
In addition, the Fortin operator $\Pi_{h}:H^{1}_{\mathrm{D}}(\Omega)^{2}\rightarrow V_{h}$ given by $\Pi_{h}w=(\tilde{\Pi}_{h}w_{n})n+(\mathcal{SZ}_{h}w_{\tau})\tau$ satisfies 42 and the stability and approximation bounds 76a, 76b and 78a stated in Assumption 3.
@@
Combining 133 and 132 shows that 76a holds.
We proceed to verify the error estimate 78a. Note that $\Pi_{h}v_{h}=v_{h}$ for any $v_{h}\in\mathbb{P}_{1}(\mathcal{T}_{h})\cap H^{1}_{\rm D}(\Omega)$. To see this, note that $\pi_{h}v_{h}=v_{h}|_{\Gamma_{\rm{T}}}$ since $v_{h}|_{\Gamma_{\mathrm{T}}}\in W_{h}$ and $\mathcal{SZ}_{h}v_{h}(z)=v_{h}(z)$ for all $z\in\mathcal{N}_{h}$ [46]. Hence,
@@
where we used the just proven stability property 76a of $\Pi_{h}$. The estimate on the $L^{2}$ error stated in 78a can now be concluded from the approximation properties of $\mathcal{SZ}_{h}$. To show 76b, one uses an inverse estimate followed by approximation properties. Namely, we have
@@
where we used the stability of $\mathcal{SZ}_{h}$ in the $H^{1}(\Omega)$-seminorm, see [8, Corollary 4.8.15]. The final error estimate for $\tilde{\Pi}_{h}$ in the $H^{1}(\Omega)$-seminorm is proven similarly. ∎
In order to check Assumption 2, we observe that the set $K_{h}$ 74 satisfies
@@
We now define ${\alpha}_{z,i}$. Let $(\hat{w}_{j},\hat{x}_{j})$, $j=0,1$, be the non-negative weights and points given by the Gauss–Radau quadrature rule [24, Table 6.1], which is an exact quadrature rule for polynomials of degree $2$ on the reference facet $\hat{E}$. Let $F_{E}$ denote the affine linear transformation from the reference facet $\hat{E}$ to a facet $E\subset\Gamma_{\mathrm{T}}$. Observe that these quadrature points define weights and points $(w_{j},x_{j})$ with $x_{j}\in\tilde{\omega}_{z}$ for $j=0,1,2,3$, such that for any continuous piecewise second-order polynomial $v$ on $\tilde{\omega}_{z}=E_{1}\cup E_{2}$, we can write
@@
Further, Assumption 5 guarantees that the weights corresponding to the boundary nodes in $\partial\Gamma_{\mathrm{T}}$ are zero in 140; i.e., $\alpha_{z}=0$ if $z\in\partial\Gamma_{\mathrm{T}}$. In turn, for each $z\in\mathcal{N}_{h}\cap\Gamma_{\mathrm{T}}$, there exists an element $T_{z}\in\mathcal{T}_{h}$, $T_{z}\subset\cup_{z_{i}}\omega_{z_{i}}$, such that if $p\in\mathbb{P}_{1}(\Delta_{T_{z}})$, it holds that
@@
(142) $\|w-\mathcal{C}_{h}w\|_{L^{2}(\Omega)}+h\|\nabla(w-\mathcal{C}_{h}w)\|_{L^{2}(\Omega)}$ $\lesssim h^{1+s}|w|_{H^{1+s}(\Omega)}$  $~\text{ for all }w\in H^{1+s}(\Omega).$
@@
We first consider the error $(\mathcal{C}_{h}v_{h}-v_{h})$ for any $v_{h}\in V_{h}$. For a node $z\in\Gamma_{\mathrm{T}}$, we consider the element $T_{z}$ such $\mathcal{C}_{h}p(z)=p(z)$ for $p\in\mathbb{P}_{1}(\Delta_{T_{z}})$; cf. 141. Then,
@@
Using that $\|\varphi_{z}\|_{L^{2}(T)}\lesssim h_{T}$, 145, and applying Cauchy–Schwarz inequality, we obtain that
@@
Let $u^{*}$ solve 124 and let $(u_{h}^{\ell},\psi_{h}^{\ell})$ come from Algorithm 3. Assume that $u^{*}\in H^{1+s}(\Omega)^{2}$, $g\in H^{1+s}(\Omega)$ and $\mathrm{div}(\mathsf{C}\epsilon(u))-f\in H^{-1+r}(\Omega)^{2}$ for some $r,s\in(0,1]$, $\|\psi_{h}^{0}\|_{L^{\infty}(\Gamma_{\rm T})}\lesssim 1$, and Assumption 5 holds. The following error estimate holds:
@@
(150) $(\mathcal{E}_{h}w)_{n}=\mathcal{C}_{h}(w_{n}-g)+g,\quad(\mathcal{E}_{h}w)_{\tau}=\mathcal{SZ}_{h}w_{\tau},$
To check Assumption 2, we first note that $(\mathcal{E}_{h}w)_{n}=0$ on $\Gamma_{\rm D}$, $(\mathcal{E}_{h}w)_{n}\in H_{\mathrm{D}}^{1}(\Omega)$ and $(\mathcal{E}_{h}w)_{\tau}\in H_{\mathrm{D}}^{1}(\Omega)$ from the properties of $\mathcal{SZ}_{h}$. This implies that $\mathcal{E}_{h}w\in H^{1}_{\mathrm{D}}(\Omega)^{2}$. We now check that $g-\mathcal{E}_{h}w\cdot n\geq 0$ on $\Gamma_{\mathrm{T}}$ to ensure that $\mathcal{E}_{h}w\in K$. For $w\in K\cup K_{h}$, we have that $(u_{h}\cdot n-g,\varphi_{z})_{\tilde{\omega}_{z}}\leq 0$ for any $z\in\Gamma_{\mathrm{T}}$ since the support of $\varphi_{z}$ on $\Gamma_{\mathrm{D}}$ is $\tilde{\omega}_{z}$. The additional Assumption 5 on the mesh guarantees that $\tilde{\omega}_{z}$ for $z\in\partial\Gamma_{\rm T}$ is not included in the definition of $\mathcal{C}_{h}$ 137. From 137, we now obtain that $\mathcal{C}_{h}(w_{n}-g)(z)\leq 0$ for all $z\in\Gamma_{\mathrm{T}}$. In addition, for the nodes on $\partial\Gamma_{\rm T}$, we find $\mathcal{C}_{h}(w_{n}-g)=-g|_{\Gamma_{\rm D}}=-\delta{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\leq}0$. Since $g-\mathcal{E}_{h}w\cdot n=-\mathcal{C}_{h}(w_{n}-g)$, we conclude that $(g-\mathcal{E}_{h}w\cdot n)\geq 0$ on $\Gamma_{\mathrm{T}}$. This verifies Assumption 2.
Noting that $\mathcal{E}_{h}w=(\mathcal{C}_{h}w_{n},\mathcal{SZ}_{h}w_{\tau})+(g-\mathcal{C}_{h}g,0)$, Assumption 3 is verified by applying Lemma 5.1, Lemma 5.2, standard estimates on $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\varepsilon=(g-\mathcal{C}_{h}g,0)$} similar to 122, and the approximation properties of the Scott–Zhang interpolant [46]. Details are skipped for brevity. ∎
@@ Acknowledgments
We thank Ioannis P. A. Papadopoulos for his comments on the manuscript and Thomas M. Surowiec and Dohyun Kim for numerous interesting discussions. BK and RM were supported in part by the U.S. Department of Energy Office of Science, Early Career Research Program under Award Number DE-SC0024335 and by the Center for Information Geometric Mechanics (CIGMO), a PSAAP-IV Focused Investigatory Center funded by the Department of Energy under Award Number DE-NA0004261. Keith’s work was also supported in part by the Alfred P. Sloan Foundation via a Sloan Research Fellowship in Mathematics. MZ acknowledges support from an ETH Postdoctoral Fellowship for the project “Reliable, Efficient, and Scalable Methods for Scientific Machine Learning”.
@@ Appendix A A stability result
The generalized Shannon entropy 30 does not fit into the form 151 for all obstacles $\phi\in L^{\infty}(\Omega_{d};\mathbb{R}^{m})$. However, the decomposition 151 is written with sufficient generality to accommodate the settings analyzed in this paper. In particular, consider the obstacle problem from Section 4, where $\phi\in H^{1}(\Omega)$ and $\phi|_{\partial\Omega}=-\delta$ for a positive constant $\delta$. In this case, 151 holds with $\mathcal{R}_{0}^{*}(\psi)=\exp(\psi)-\delta\psi$, $\phi_{0}=1$, and $\phi_{1}=\phi+\delta$, implying $\nu_{1}=\delta$. For the Fermi–Dirac entropy given in Example 2.2, which is a suitable choice for bilateral obstacle problems ($B=\mathrm{id}$ and $Q=V=H^{1}_{0}(\Omega)$), we can write
@@
Thus, defining $\mathcal{R}_{0}^{*}(\psi)=2\ln(\exp(\psi)+1)-\psi$ and assuming that $\overline{u}-\underline{u}\geq\underline{\phi_{0}}$ and $\overline{u}+\underline{u}\in H^{1}_{0}(\Omega)$, we deduce that 151 holds with $\nu_{1}=1$. Notably, if $\underline{u},\overline{u}\in\mathbb{R}$ with $\underline{u}<0<\overline{u}$, the condition $\underline{u}+\overline{u}\in H^{1}_{0}(\Omega)$ is not generally satisfied. However, in this case, we can still verify 151 using an alternative decomposition given by $\mathcal{R}_{0}^{*}(\psi)=(\overline{u}-\underline{u})\ln(\exp(\psi)+1)+\underline{u}\psi$, $\phi_{0}=1$, and $\phi_{1}=0$. The Hellinger entropy, introduced in Example 2.3, readily satisfies 151 with $\mathcal{R}_{0}^{*}(\psi)=\gamma\sqrt{1+|\psi|^{2}}$, $\phi_{0}=1$, and $\phi_{1}=0$, implying $\nu_{1}=\gamma$ and $c_{1}=0$.
@@
Assume that 151 holds. Further, assume that $\psi_{h}^{0}$ is selected so that $\|B^{\prime}\psi_{h}^{0}\|_{V^{\prime}}\lesssim 1$. Then there exists a constant $C_{\mathrm{stab}}$, independent of $h$, $\ell$, and $\{\alpha_{k}\}_{k=2,\ldots}$, such that
@@
We first observe that testing 10a with $v_{h}=u_{h}^{k}$ and 10b with $w_{h}=\psi_{h}^{k}$, subtracting the resulting equations, and using the expression 151a yields
@@
Noting that $c_{1}\leq\mathcal{R}_{0}^{*}(0)$ by 151b, the above estimates yield
@@
Employing the coercivity of the bilinear form $a$ 8, we obtain that
@@
where $\beta>0$ is the same constant appearing in the LBB condition 41. Using the above bound in 157 and appropriate applications of Young’s inequality shows that there exists a constant $M_{k}$ (depending only on $\alpha_{k}$, $M$, $\|F\|_{V^{\prime}}$, $\|\Pi_{h}\|$, $\|\phi_{1}\|_{Q}$, $\beta$, and $c_{2}$) such that
@@
Using 8, we obtain that
@@
where we used (10a) and similar arguments to (159). The bounds 164 and 165 yield the first two terms in 153.
@@
where $C_{3}=(\|F\|_{V^{\prime}}+2\nu^{-1/2}MC_{2}+\|B^{\prime}\psi_{h}^{0}\|_{V^{\prime}})\|\Pi_{h}\|$. To derive the bound on $\|\psi_{h}^{k}\|_{L^{1}(\Omega_{d})}$, we substitute 168 into 157. ∎